# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

### v2

@@ 2.4 Training
The training, validation, and testing sets were randomly split in a ratio to sizes . During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle. Cross-entropy loss was used, weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C. The model from the training epoch that maximized balanced validation accuracy was chosen as the final model. Both CRNN and RNN models were trained with Adam [32], a widely-used adaptive learning rate optimizer.
The CRNN model was trained on single batches. Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation. The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
Three models of varying sleep stage resolution were built: a 3-class, 4-class, and 5-class sleep staging model. The 5-class model has classes for wake, N1, N2, N3, and REM sleep, while the 3-class model groups N1, N2, and N3 sleep into non-REM sleep. The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [4].
@@ 2.5 Ensembling
For each of the classes inside a -class ensemble model, a binary model was trained to predict the presence or absence of that class. The logits of these binary models were stacked with logits from the regular -class models to form the inputs to a suite of decision tree-based architectures including AdaBoost [33], random forests [34], extra-tree classifiers [35], histogram gradient boosting [36], and XGBoost [37]. The model among the suite that maximized balanced validation accuracy was selected as the final ensembler model. A visual representation of the ensembling pipeline is presented in Figure 2C.
@@ 2.6 Statistical Analyses
Results were compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models. Overall performance was evaluated by aggregating all test set recordings and computing the balanced accuracy, weighted precision, weighted recall, weighted F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC).
For the best model type, performance was plotted on a healthy subset of the test set and ablation studies are carried out using various subsets of the ANNE One sensor data and a different sequential architecture. For the best 5-class model, Blant-Altman plots were created to evaluate class-wise biases, followed by analyses of averaged per-recording metrics both individually and across clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency. Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
@@ 3.1 Model Evaluation
Typical-case recording predictions for all 5-class models are shown in Figure 4 along with predicted point-wise output class probabilities for models that can produce them. A Uniform Manifold Approximation and Projection (UMAP) [38] of an intermediate CRNN output is also provided in Appendix E, which shows the separation of sleep stages as clusters in the learned representation space.
Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. No meaningful deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models), of which the latter is presented in this paper. The RNN ensemble model yields the best metrics and dominates most class-wise metrics (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, Appendix F). Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count. Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
[…]
In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model experiences a significant boost in performance as denoted in Figure 6 although removal of these participants from the test set does not substantially worsen model performance, indicating that overall model performance is not driven by these healthy participants (Appendix G, Figure 12).
@@ 3.2 Feature Ablation Study
We next investigated which sets of sensors (chest module vs. limb module; ECG vs. PPG. vs. other) were most influential to model performance. To do so, modified instances of the RNN ensemble model were trained to determine the utility of its components. Firstly, feature sets solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined. As shown in Table 2, the chest-only model trails closely behind the original in performance while the limb-only model exhibits a marked degradation () for all metrics. This strongly suggests that the chest module is the principal driver behind model accuracy, although the limb module alone is capable of recovering sleep staging information to a rough degree. Next, because the ECG and and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of RNN ensemble models trained on only the accelerometry and temperature sensors. Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals.
[…]
3.3 Recurrent Architectures
In order to understand whether the incorporation of Mamba improved the RNN ensemble model, a similarly-sized variant model replacing Mamba with LSTM was trained, its results provided by Table 3. The LSTM model shows a minor loss in all metrics, confirming the Mamba approach is beneficial.
[…]
@@ 3.4 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs. PSG actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating wake and N2 sleep and slightly overestimating N1 and N3 sleep.
3.5 Performance by Clinical Group
Averaged recording-wise metrics are catalogued in Table 4 for the 5-class RNN ensemble, wherein linear regressions were performed to fit various recording metrics to accuracy (plots in Appendix G, Figure 13). Performance of the RNN ensemble is robust to differences in all metrics except AHI, which had a weak negative correlation with accuracy, though the proportion of variance in accuracy explained by AHI is modest (RNN ). Four annotators were involved in generating the ground truth labels used in this study and the agreement between the RNN model and the ground truth labels did not differ by the individual who generated the annotations (RNN , one-way F-test).
@@ 3.5 Performance by Clinical Group
Periodic Limb Movement Index (PLMI) | 0.0338 | 0.1830 | ||
Apnea-Hypopnea Index (AHI) | 0.0751 | 0.0449 | ||
Body Mass Index (BMI) | 0.0350 | 0.1753 | ||
Total Sleep Time (TST) | 0.0055 | 0.5940 | ||
@@ 4.1 Key Findings
This work has developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages. The models demonstrate good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
@@ 4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 5 & 7) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
@@
Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity, which may influence the relationship between sleep stage and its autonomic, cardiac, and pulmonary manifestations. In contrast, we studied an older (mean age 57.49) set of participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6). Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG.
@@
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the best person-specific model in the paper which had a 4-class balanced accuracy of 73.4%.
[…]

### v4

@@ 2.4 Training
The training, validation, and testing sets were randomly split in a ratio to sizes . During the training process, recordings were clipped to random segments of size 5000 to allow for mini-batch training () and to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle. The objective function was cross-entropy loss, weighted on inverse class frequency to adjust for the moderate class imbalance shown in Figure 1C, then averaged across batches and time windows. The model was optimized with Adam [kingma], a widely-used adaptive learning rate optimizer, with learning rate 0.001. The parameters from the training iteration that maximized balanced validation accuracy were chosen for the final model.
Three models of varying sleep stage resolution were built: a 3-class, 4-class, and 5-class sleep staging model. The 5-class model has classes for wake, N1, N2, N3, and REM sleep, while the 3-class model groups N1, N2, and N3 sleep into non-REM sleep. The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [imtiaz2021systematic]. Separate models were trained for each class count as to treat each class equally.
@@ 2.5 Ensembling
For each of the classes inside a -class ensemble model, a new binary model of the same architecture was trained from scratch to predict the presence or absence of that class. The logits of these binary models were then stacked with logits from the regular 3/4/5-class models. This collection of logits formed the inputs to a suite of decision tree-based architectures (see S.3 for details) including AdaBoost [freund1995desicion, scikit-learn], random forests [ho1998random, scikit-learn], extra-tree classifiers [geurts2006extremely, scikit-learn], histogram gradient boosting [cui2021gbht, scikit-learn], and XGBoost [chen2016xgboost]. This architecture type processes time windows in a time-independent manner, progressively dividing them into smaller and more homogeneous groups by thresholding based on the value of their features (i.e. a logit value for a sleep stage); the thresholds and choice of features are learned parameters. Sleep stage predictions in a single decision tree are made from the plurality sleep label of training set time windows inside a terminal node, while the proportions of each sleep label become the class probabilities. The model among the suite that maximized balanced validation accuracy was selected as the final ensembler model. A visual representation of the ensembling pipeline is presented in Figure 2C.
@@ 2.6 Statistical Analyses
Results were compiled of 3-class, 4-class, and 5-class sleep staging for the regular and ensemble RNN models. Overall performance was evaluated by aggregating all test set recordings and computing the balanced accuracy, weighted precision, weighted recall, weighted F1, Cohen’s kappa (), and Matthews correlation coefficient (MCC).
For the best model, performance was plotted on a healthy subset of the test set. Ablation studies were carried out using various subsets of the ANNE One sensor data. We also retrained the model without downsampling the data, and considered alternative architectures including use of LSTM instead of Mamba, and consideration of a CRNN model utilizing the same preprocessing steps as our RNN model. For the best 5-class model, Bland-Altman plots were created to evaluate class-wise biases, followed by analyses of per-recording metrics individually as well as across differing signal quality scenarios, sleep disorders, and clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency, and sleep disorders including insomnia, restless legs syndrome, and REM sleep behaviour disorder. Because individual recordings did not necessarily contain instances of all five sleep stages, balanced accuracy may not have been a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
@@ 3.1 Model Evaluation
Typical-case recording predictions for all 5-class models are shown in Figure 3 as hypnograms along with predicted point-wise output class probabilities, known as hypnodensity graphs [anderer]. A Uniform Manifold Approximation and Projection (UMAP) [umap] of an intermediate model output is also provided in S.4 (Figure S4), which is able to show the separation of sleep stages as clusters in the learned high-dimensional representation space.
Figure 4 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. We plotted performance in both 1-second and 30-second epochs and found similar results irrespective of the epoch length (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models). Accordingly, all results in this paper are for 30-second windows. Compared to a base RNN model, the ensembling procedure resulted in improved overall performance (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, S.5). Most misclassifications were made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
![[Uncaptioned image]](figures/ensemble_outputs_rnn.png)
Metric (%) Classes Model Balanced Accuracy Precision Recall F1 Cohen’s MCC RNN 60.20 69.48 61.20 63.68 49.16 50.18 5 RNN Ensemble 65.11 71.61 64.41 66.15 53.23 54.38 [1pt/1pt] RNN 67.59 73.08 66.59 68.34 51.95 53.06 4 RNN Ensemble 75.30 75.80 73.68 74.10 61.51 61.95 [1pt/1pt] RNN 81.09 83.20 81.98 82.35 69.59 69.77 3 RNN Ensemble 84.02 84.76 84.04 84.23 72.89 73.00
In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model performed even better as denoted in Figure 5 although removal of these participants from the test set did not substantially worsen model performance (S.7, Figure S7), indicating that overall model performance was not driven by this small number of healthy participants.
![[Uncaptioned image]](figures/healthy-subset.png)
@@ 3.2 Feature Ablation Study
We next investigated which sets of sensors (chest module vs. limb module; ECG vs. PPG vs. other) were most influential to model performance. Modified instances of the ensemble model were trained to determine the utility of its components. Firstly, features solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined. As shown in Table 1, the chest-only model trails closely behind the original while the limb-only model exhibits a marked degradation () for all metrics. This strongly suggests that the chest module is the principal driver behind model accuracy, although the limb module alone is capable of recovering sleep staging information to a rough degree. Next, because the ECG and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of a model trained on only the accelerometry and temperature sensors. Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals, with temperature providing additional complementary information. Further exclusion of temperature data resulted in an accelerometry-only model that performed similarly to or even slightly better (Table S3), indicating that chest accelerometry alone captures most of the discriminative information used by the network. This is corroborated by a per-sensor feature importance analysis with Integrated Gradients [sundararajan2017axiomatic] (Figure S17), which suggests that accelerometry-derived features contribute the most to model performance relative to temperature, PPG, and ECG-derived features.
Metric (%) Classes Feature Set Balanced Accuracy Precision Recall F1 Cohen’s MCC Full 65.11 71.61 64.41 66.15 53.23 54.38 Chest Module 63.94 70.66 62.63 64.22 51.23 52.59 Limb Module 54.83 62.29 55.08 56.56 41.60 42.59 5 No ECG/PPG 63.85 70.68 63.35 65.03 51.97 53.12 [1pt/1pt] Full 75.30 75.80 73.68 74.10 61.51 61.95 Chest Module 73.46 74.63 71.64 72.26 58.88 59.47 Limb Module 64.89 68.63 63.59 64.66 48.22 49.13 4 No ECG/PPG 73.03 74.46 71.52 72.10 58.70 59.30 [1pt/1pt] Full 84.02 84.76 84.04 84.23 72.89 73.00 Chest Module 81.77 83.58 82.50 82.83 70.39 70.53 Limb Module 73.28 78.10 74.71 75.84 58.21 58.69 3 No ECG/PPG 82.78 84.26 83.18 83.51 71.71 71.56
To better characterize the information available in chest accelerometry, we probed the signals within these channels. At the individual-participant level, aggregated accelerometry spectrograms aligned to the hypnogram (Section S.8) show stage-dependent structure, with clear shifts in power concentrated in a low-frequency band overlapping with the range of expected respiratory rates in sleep. This supports the hypothesis that very low-frequency components of chest accelerometry are important for distinguishing sleep stages. To examine this quantitatively, we (1) trained a model using accelerometry channels alone and then (2) replaced the band-pass filter with a high-pass cutoff to Hz to attenuate the expected frequency range of resting respirations. Results are shown in Table S3. The accelerometry-only model performed marginally better than the No ECG/PPG model but worse than the full model; this may be attributable to the data’s unimodal nature, which could reduce the complexity of learning structural patterns in multiple data types during training, thus improving convergence. Meanwhile, the reduced performance of the high-passed accelerometry-only model is compatible with the possibility that some information used by the model is in the frequency range of human respirations.
3.3 Alternative Approaches
To guard against overfitting and reduce the computation burden, we downsampled the ECG, PPG, and accelerometry signals to 100Hz prior to use in our model. To explore the extent to which this would have impacted performance, we retrained the model without downsamping (i.e. using 512 Hz ECG, 128 Hz PPG, and 210 Hz accelerometry data) and band-pass filtering. This did not result in any decisive improvement in model performance (S.7, Figure S6). A strong correlation between high-frequency and low-frequency bins might have contributed to redundancy and increased the likelihood of overfitting to noise. In support of this hypothesis, we note the difference in balanced accuracy between the training and test sets was much less in the downsampled data () than in the non-downsampled data ().
Two variants of the RNN model were produced to determine the value of the Mamba architecture and the 1-second window approach. The first, a LSTM-based RNN, replaces the Mamba blocks with LSTM blocks of identical input, hidden, and output dimension sizes, resulting in a comparable parameter count to the original RNN model. Following our previous study [sleep_poster], the second is a CRNN (1271K parameters) that groups every 30 time windows into buckets, extracts bucket features with a CNN, and processes features sequentially using a Mamba-based RNN, resulting in predictions made per 30 seconds (see S.6 for details). The technique of extracting features from 30 seconds of data for an autoregressive model has seen many variants [sun, sleepppg, zhang2018sleep], making the CRNN suitable as a "baseline" approach to windowing. As provided by Table 2 (confusion matrices in S.7, Figure S6), both the LSTM-based RNN models and the CRNN models have modestly reduced performance in all metrics, confirming the Mamba approach and 1-second window approach are beneficial.
Metric (%) Classes Model Balanced Accuracy Precision Recall F1 Cohen’s MCC CRNN 60.02 56.30 60.02 56.37 46.96 47.92 5 LSTM-Based RNN 62.64 69.46 62.47 64.03 50.76 51.87 RNN 65.11 71.61 64.41 66.15 53.23 54.38 [1pt/1pt] CRNN 70.59 63.68 70.59 66.09 55.13 55.47 4 LSTM-Based RNN 72.82 73.95 71.64 72.06 58.70 59.17 RNN 75.30 75.80 73.68 74.10 61.51 61.95 [1pt/1pt] CRNN 80.21 73.43 80.21 75.76 65.26 65.46 3 LSTM-Based RNN 81.06 82.90 81.37 81.82 68.71 68.95 RNN 84.02 84.76 84.04 84.23 72.89 73.00
@@ 3.4 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 6 contains scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing predicted vs. actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating N2 sleep and wake (as shown by the overestimated sleep in the first column) and slightly overestimating N1 and N3 sleep.
3.5 Recording-Wise Performance
We computed recording-wise performance metrics for the 5-class RNN ensemble model, the distributions of which are summarized in Table 3 . We then considered recording-level accuracy as a function of clinical and demographic features, and of ECG and PPG signal quality, using linear regression models or -tests (plots in S.7, Figures S8, S9, S11). The model performed similarly on recordings from individuals with insomnia as compared to those without (accuracy for those with as compared to without insomnia, ) and in those with RLS compared to those without (accuracy , ).
Accuracy sunk for those with REM sleep behaviour disorder (accuracy , ), largely driven by poorer classification of REM sleep (39% and 79% of REM sleep was detected by the 5-class RNN ensemble for those with and without RBD respectively). Accuracy also declined with higher AHI (); however, this was a modest effect, with accuracy dropping only 2% for every 10 unit higher AHI. Accuracy was not significantly associated with age, sex, periodic limb movement index, sleep efficiency, or total sleep time.
Model accuracy did not vary significantly with the proportion of the recording with bad quality ECG (estimate % decline in accuracy for each 1% increase in the proportion of poor-quality ECG, ) or PPG (estimate % decline in accuracy per 1% increase in the proportion of poor-quality PPG, ), but it was significantly worse in recordings with a high proportion of epochs with neither good ECG nor good PPG SQI (estimate % decline in accuracy for each 1% increase in the proportion of time where both ECG and PPG SQI were poor, ). However, this effect was modest () and we note that the proportion of time where this occurred was generally low, with all but 6 recordings having at least one of ECG or PPG with good quality for at least 90% of the recording. Even in recordings in which 40% or more of the recording had neither good quality ECG nor good quality PPG, 5-class model accuracy remained above 0.50.
Four annotators were involved in generating the ground truth labels used in this study and the agreement between the model and the ground truth labels did not differ by the individual who generated the annotations (, one-way F-test).
@@ 3.5 Recording-Wise Performance
| Periodic Limb Movement Index | 0.0338 | 0.1830 | ||
| Apnea-Hypopnea Index | 0.0751 | 0.0449 | ||
| Body Mass Index | 0.0350 | 0.1753 | ||
| Total Sleep Time | 0.0055 | 0.5940 | ||
@@
| Disorder | ||||
| Restless Leg Syndrome (9) | 0.0134 | 0.4044 | ||
| Insomnia (14) | 0.0135 | 0.4032 | ||
| REM Sleep Behaviour Disorder (5) | 0.0965 | 0.0222 | ||
| % of Recording with Good SQI In… | ||||
| ECG | 0.0623 | 0.0687 | ||
| PPG | 0.0005 | 0.8753 | ||
| ECG or PPG | 0.0913 | 0.0264 | ||
| ECG and PPG | 0.0136 | 0.4017 | ||
@@ 4.1 Key Findings
This work has developed a Mamba-based RNN ensemble model to accurately perform inference from the ANNE One device for major sleep stages. The model demonstrates good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality.
@@ 4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 4 & 5) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [radha2019sleep, sun, sleepppg, pini, topalidis2023pulses, jones2024expert, kazemi2024improved] although a handful of studies used data from purpose-designed wearable sensors [zhang2018sleep, sun, anne_personal, pini, topalidis2023pulses, silva, wulterkens:2021]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the macro performance of our ensemble RNN model approached [sun, topalidis2023pulses, jones2024expert, radha2019sleep, sleepppg] or exceeded [pini, kazemi2024improved] the performance reported for models trained on much larger sets of PSG data.
@@
Our model exceeded [silva, pini, anne_personal, sun, zhang2018sleep] or matched [wulterkens:2021, schipper] the benchmarks of most wearable sensor based models with the exception of one [topalidis2023pulses]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity, which may influence the relationship between sleep stage and its autonomic, cardiac, and pulmonary manifestations. In contrast, we studied an older (mean age 57.49) set of participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 5). Another contrast between [topalidis2023pulses] and the present study is that [topalidis2023pulses] performed much stricter QC for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG, such that sensors returning poor-quality data are compensated for in the model by those returning good quality data. Indeed, poor signal quality was present in at least one sensor module for 866 (29.7%) of the 2914 hours of recording in our dataset. However, for 725 (83.7%) of these hours, the other module returned sufficient quality signal to allow accurate predictions.
@@
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [anne_personal]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the best person-specific model in the paper which had a 4-class balanced accuracy of 73.4%.
Study (Year) Dataset Metric % Signal Features Radha et al. (2019) [radha2019sleep] SIESTA (7 sleep labratories) 584 recordings (292 subjects) 25% testing 4-stage Acc. 4-stage 77 61 ECG [1pt/1pt] Sun et al. (2020) [sun] MGH (sleep laboratory) 8682 recordings (7208 subjects) 20% testing 3-stage B.Acc. 3-stage F1 3-stage 5-stage B.Acc. 5-stage F1 5-stage 86.8 84.2 76.0 70.3 68.1 58.5 ECG, Resp. Effort [1pt/1pt] Kotzen et al. (2022) [sleepppg] SHHS Visit 1 (multi-center) 5758 recordings (pretraining) MESA (general population) 2056 recordings 204 testing 4-stage Acc. 4-stage B.Acc. 4-stage F1 4-stage 84 76 83 75 PPG [1pt/1pt] Kotzen et al. (2022) [sleepppg] SHHS Visit 1 (multi-center) 5758 recordings (pretraining) CFS Visit-5 v1 (family study) 324 recordings 80 testing 4-stage Acc. 4-stage B.Acc. 4-stage F1 4-stage 82 80 82 74 PPG [1pt/1pt] Pini et al. (2022) [pini] Unspecified (academic sleep centers) 12404 recordings CinC (sleep laboratory) 994 recordings External testing set 3-stage B.Acc. 3-stage F1 3-stage 3-stage MCC 4-stage B.Acc. 4-stage F1 4-stage 4-stage MCC 74.0 81.4 61.6 61.7 63.6 72.0 53.6 54.0 ECG [1pt/1pt] Topalidis et al. (2023) [topalidis2023pulses] Unspecified (general population w/ sleep complaints) 314 recordings (185 subjects) 25% testing Ambulatory home recordings 4-stage B.Acc 4-stage F1 4-stage 4-stage MCC 85.4 86.3 79.2 79.2 ECG [1pt/1pt] Jones et al. (2024) [jones2024expert] CCSHS, CFS, CHAT, MESA, WSC 4000 random sampled recordings 500 testing 5-stage B.Acc 5-stage Median 74 72.5 ECG [1pt/1pt] Kazemi et al. (2024) [kazemi2024improved] UC Irvine Sleep Center 123 recordings 20% testing 3-stage 4-stage 5-stage 71.4 55.0 61.6 PPG, Resp. Flow Rate, Resp. Effort
Study (Year) Dataset Metric % Signal Features Zhang et al. (2018) [zhang2018sleep] Beijing General Hospital of the Air Force (general population) 39 recordings 12.5% testing 5-stage Prec. 5-stage Rec. 5-stage F1 58.5 61.1 58.5 Heart Rate, Actigraphy [1pt/1pt] Sun et al. (2020) [sun] Model from Table 4 SHHS Visit 1 & 2 (multi-center) 1000 random sampled recordings External testing set 3-stage B.Acc. 3-stage F1 3-stage 5-stage B.Acc. 5-stage F1 5-stage 80.8 80.2 69.7 63.9 58.6 53.3 ECG, Resp. Effort [1pt/1pt] Wulterkens et al. (2021) [wulterkens:2021] SOMNIA, N2N, HHS, HealthBed 835 recordings 35% testing 4-stage B.Acc. 4-stage F1. 4-stage 4-stage MCC 75.4 76.4 63.7 63.7 PPG, Accel. [1pt/1pt] Chen et al. (2022) [anne_personal] Individual-Based Model Shirley Ryan AbilityLab (inpatient unit) 10 recordings 10% testing 3-class Mean 4-class B.Acc. 4-class Mean 60.0 73.0 61.7 ECG, PPG, Accel., Temp. [1pt/1pt] Chen et al. (2022) [anne_personal] Population Model Shirley Ryan AbilityLab (inpatient unit) 10 recordings 10% testing 3-class Mean 4-class Mean 17.1 6.1 ECG, PPG, Accel., Temp. [1pt/1pt] Pini et al. (2022) [pini] Model from Table 4 Z3Pulse (general population) 156 recordings (52 subjects) ECG-based wearable External testing set 3-stage B.Acc. 3-stage F1 3-stage 3-stage MCC 4-stage B.Acc. 4-stage F1 4-stage 4-stage MCC 72.5 79.8 61.2 61.4 64.2 69.8 53.2 53.6 ECG [1pt/1pt] Topalidis et al. (2023) [topalidis2023pulses] Model from Table 4 Unspecified (general population w/ sleep complaints) 314 recordings (185 subjects) 25% testing 4-stage B.Acc 4-stage F1 4-stage 4-stage MCC 84.4 83.9 76.2 76.3 ECG, PPG [1pt/1pt] Schipper et al. (2024) [schipper] Sleep Medicine Center Kempenhaeghe, OLVG Hospital (at-home recordings) 323 recordings 4-stage B.Acc. 4-stage F1 4-stage 4-stage MCC 75.6 79.4 67.1 67.2 Accel. [1pt/1pt] Silva et al. (2024) [silva] Instituto do Sono (general population) 1522 recordings (1430 subjects) 586 testing 4-stage Mean Acc. 4-stage Mean 70.7 56 Heart Rate, Accel.